# BAP v2: An Enhanced Task Framework for Instruction Following in Minecraft Dialogues

### v1

@@ 1 Introduction
The MCBT situates a dialogue task—designed to be asynchronous, asymmetric, and with minimal language constraints—in a simulated 3D Minecraft environment where participants, with constantly shifting perspectives, must refer to and modify a dynamic world. In this task, two players collaborate: an Architect (A) instructs a Builder (B) to construct a target structure using multi-colored blocks (Figure 1). The MDC comprises 509 human-human game logs for this task. Narayan-Chen, Jayannavar, and Hockenmaier (2019) focus on generating Architect utterances, and Jayannavar, Narayan-Chen, and Hockenmaier (2020) focus on end-to-end neural models for building an automated Builder agent. The latter defined the challenging Builder Action Prediction (BAP) task, aimed at predicting the sequence of actions (block placements and/or removals) a human Builder performed at a specific point in a human-human game. Despite the task’s inherent challenges and limited training data, we demonstrate that models trained with a suitable amount of game history, enriched world representations, and a sufficient amount of synthetic, diversified data (via simple data augmentation techniques) yield promising initial results. We recap relevant prior work on the Minecraft Builder in Section 2. Building on this foundation, we next take a closer look at evaluation and training data for the BAP task, uncovering key challenges and making significant improvements on both fronts to propose BAP v2, an upgraded version of the task. This will allow future work to make more efficient and meaningful progress on it.
@@
Our work on evaluation highlights spatial reasoning as a key bottleneck in model performance, alongside the challenge of limited training data (even with the aforementioned data augmentation). Instead of merely collecting more data (which can be impractical/expensive in realistic scenarios for such complex situated dialogue tasks), we create additional diverse synthetic training data generated from novel Minecraft dialogue and target structure simulators that emulate the MCBT (Section 5). This data, though naturally simpler than it, is rich in spatial relations and referential expressions and we show that it can be successfully used to train better models, even with simple training methods (Section 6). We demonstrate that jointly training on combined synthetic and original BAP (augmented) data already improves performance. Further, a straightforward Curriculum Learning method provides an additional boost. Together, the synthetic and augmented BAP data form the BAP v2 training set. Looking ahead, such data could also be crucial for training more sophisticated, data-hungry deep transformer models and training/fine-tuning increasingly large LLMs. We provide a detailed analysis of model performance of our best performing model relative to the baseline in Section 7.
@@
Section 8 provides a detailed discussion on four topics – an additional key challenge in BAP evaluation, the importance of evaluation on the synthetic data, relevant concurrent work, and broader implications of our work that go beyond BAP. The latter further highlights the potential impact of our work and directions for future research. Related work is covered in Section 9. We conclude and discuss future work in Section 10.
@@ 2.1.2 MCBT subtasks: Architect Utterance Generation and Builder Action Prediction
We define the Architect Utterance Generation (AUG) Task as the task of generating a suitable Architect utterance at any point in a human-human game at which the human Architect produced an utterance (Narayan-Chen, Jayannavar, and Hockenmaier, 2019), and the Builder Action Prediction (BAP) Task as the corresponding task of generating a suitable Builder action sequence (consisting of block placements/removals) at any point in a human-human game at which the human Builder placed or removed blocks (Jayannavar, Narayan-Chen, and Hockenmaier, 2020). Although both of these tasks consider only a single turn, and assume the context of a game and dialogue between two human players, they are important first steps towards the creation of agents that can complete an entire game in the MCBT with a human counterpart.
@@ 2.5 Data Augmentation
Since the small size of the training set (3,709 examples) is a major limiting factor for training complex models, in Jayannavar, Narayan-Chen, and Hockenmaier (2020), we generated synthetic variants of the original game logs in the training data by combining three data augmentation techniques: utterance paraphrasing (synonym-based substitutions), color substitution (random color permutations across logs), and spatial transformations (rotating structures and Builder B) (details in appendix). Empirically, we found that increasing the training data to 14,836 (4x) items gave the best performance for our baseline model (Section 2.6). We will refer to this as the original BAP (augmented) data.
@@ 2.6 A CNN- and GRU-based baseline model
In Jayannavar, Narayan-Chen, and Hockenmaier (2020), we proposed a neural model for the BAP task. In this work, we build upon it and refer to it as the baseline model. The model (Figure 4) is based on a recurrent encoder-decoder architecture (Sutskever, Vinyals, and Le, 2014; Cho et al., 2014) in which a GRU-based encoder (bottom left box) captures the game context (dialogue and action history) via GloVe embeddings (Pennington, Socher, and Manning, 2014), and a CNN-based encoder (top left box) captures the world state at each time step. The world state encompasses the current built structure, action history and the Builder’s position and orientation. The decoder (right box) predicts one action per time step, based on the game history, the world state at that time, and the last action taken. It consists of another GRU backbone over action sequences (bottom right), and a multi-class classifier that reads in the output of the GRU backbone as well as the world state encoding produced by the CNN to predict either the next action (block placement or removal) to be taken, or a special STOP token that terminates the action sequence. Actions are represented as feature vectors and embedded via feedforward layers before being fed to the decoder GRU. The world state representation gets updated and re-encoded after each predicted action. We train this model with teacher forcing and cross entropy loss on the original BAP (augmented) data, and use greedy decoding (max. sequence length of 10 actions) at test time. This yields an F1 score of 21.2% on the BAP test set. More details about the model are provided in the appendix.
@@ 3.1 Annotating the test set
We manually annotate items in the test set so that we can identify cases that have to be distinguished for evaluation purposes. Annotators were asked to annotate all items in the same dialog in sequence. For each non-empty board (NEB) item, annotators were first asked whether the context is clear and has a unique interpretation , i.e. clearly specifies what structure to be built next, and there is only a single way of placing that structure on the board (as is typically the case when the board is not empty), or whether the context is unclear, i.e. it is unclear what the Builder should do next, either because more information is required, or no new instruction was given. For empty board (EB) items, annotators could also indicate that the context is clear and has multiple interpretations , i.e. it is clear what structure should be built, even though there are multiple possible placements or orientations within the build region, as is often the case when the board is empty. If the context is clear (with unique or multiple interpretations), a second question asked whether the built structure is correct or incorrect. Details about the annotation interface, the annotation workflow, the annotation tool, data collection setup, etc. are provided in the appendix. We now illustrate the resulting categories of BAP items with a few examples.
@@ 5.1.2 High-level Algorithm
    Optional Clarification Exchange: Some information may be omitted from the instruction with a small probability, allowing B to ask a clarification question. An appropriate exchange is then generated, with B asking and A providing the missing information. E.g., in Figure 11, the block’s color is omitted, prompting B to ask, "what color?" and A to reply, "red".
@@
We implement this general framework in three distinct ways, resulting in three different simulators. Each simulator uniquely instantiates steps 1, 3, and 4, along with independent target structure generation, while steps 2, 5, and 6 remain largely consistent across simulators. For brevity, we provide only a summary of each simulator in the following sections, with detailed descriptions in the appendix.
@@ 5.2 Random target structures and dialogs
The dialogs correspond to random target structures generated dynamically during the dialog simulation (Section 5.1.2) which runs for a maximum number of iterations. Blocks are placed/removed randomly, and one at a time in each iteration, with limited planning on A ’s part for block placement. As a result, the target structure is disordered, and A ’s instructions are fairly low-level, addressing only one block at a time. Most of them use spatial relations relative to arbitrary blocks identified by color and location (e.g., "to the right of the red" or "to the right of the bottom-most red block"). Figure 12 shows an example BAP item from this dataset. A identifies the orange block as a reference block and uses a 1D spatial relation "behind" to specify the target yellow block.
@@ 5.3.1 Shape-based structures
We define six unique elementary 3D shapes: rows, diagonals, T-shapes, L-shapes, U-shapes, and planes. These shapes were chosen for their simplicity and occurrence in MDC dialogs and structures. Formal definitions of these shapes are provided in the appendix. We design a simulator that generates target structures by combining multiple randomly sampled elementary shapes into a single composite structure. Sampling is performed with replacement, allowing multiple instances of the same shape type within a structure. For each shape instance, properties such as size, color, orientation, and exact 3D location are randomly sampled within specific constraints (e.g., minimum and maximum size bounds). Some global constraints are also enforced (similar to those for random structures in Section 5.2, with an added constraint for shapes):
@@ 5.3.2 Blocks-based dialogs for shape-based structures
The dialog simulation (Section 5.1.2) takes the target structure as input and continues iteratively until the target structure is built. Construction proceeds one shape at a time. A ’s planning is involved and follows a heuristic that selects one or more blocks to place in each iteration until all shapes are completed. Consequently, A ’s instructions reference a single block or multiple ones collectively, with each shape decomposed into multiple instructions. Most of A ’s instructions include spatial relations relative to the last block placed by B (e.g., "now, place two red blocks behind the last block"). Figure 15 illustrates an example BAP item from this dataset. The first blue square has been built, and the second (yellow) shape is under construction. A is instructing B to continue building the latter by placing three more blocks, using a 1D spatial relation "on top of" relative to the last block. (In the previous step, A used a 2D spatial relation, "diagonally in front of and to the right of".)
@@ 5.3.3 Shape-based dialogs for shape-based structures
The dialog simulation (Section 5.1.2) takes the target structure as input and proceeds shape by shape, with one complete shape placed per iteration until the target is built. A ’s planning employs simple heuristics to determine the order of shapes and select a target block within each shape for defining spatial relations, which also serves as B ’s starting point for building the shape. A ’s instructions, therefore, describe entire shapes rather than individual blocks, resulting in higher-level, abstractive instructions. Since each target structure consists of two shapes, each placed in one go, a 1D/2D/3D spatial relation is only used to specify the position of the initial block of the second shape relative to the last block placed in the first shape. Additionally, A can also use spatial relations that use B as a spatial anchor (Section 5.1.3), for building either shape. Figure 16 illustrates a BAP item from this dataset. The first shape (a yellow column) has been built, and A is instructing B to build the second shape (a red row) in one go. A uses the spatial relation "going to the left of you" (using B as a spatial anchor) to indicate the direction for the row, along with the 3D spatial relation "one to the left, two underneath, and one block in front of" relative to the last block (the topmost yellow one), to indicate the starting position for the row.
Figure 16: A BAP item from the shapes-based dialogs for shape-based structures
@@ 6 Training with Synthetic Data
Our next objective is to demonstrate that the synthetic data can help train much better models for the BAP task, even with simple training methods, thereby addressing the challenge of insufficient training data. We first show that simple joint training on the aggregation of the synthetic data and the original BAP (augmented) data already yields a better model (Section 6.1). We then go a step further and show that a simple Curriculum Learning method can further boost performance (Section 6.2). The synthetic data together with the original BAP (augmented) data now constitutes the BAP v2 training dataset, completing the BAP v2 framework. We also illustrate the impact of the synthetic data and training methodologies on training a simple LLM (BERT) and transformer-based model (Section 6.3).
@@ 6.1.1 Experimental Setup
Our setup closely follows that of Section 2.6, where the baseline $M_{\mathrm{mc}}$ was trained on $D_{\mathrm{mc}}$ alone. We retain the same exact model architecture and hyperparameters as in $M_{\mathrm{mc}}$. Training employs teacher forcing with cross-entropy loss, and decoding is performed via greedy decoding, with a maximum sequence length of 20 actions. We use the fairer F1 metric to report micro F1 scores on the BAP v2 test set for $D_{\mathrm{mc}}$ (Section 4.4) and the test sets of $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$.  [^6] For fairer F1 on $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $\mathcal{M}$ consists of all EB items, unlike $D_{\mathrm{mc}}$, where only a subset is considered (Section 4.1). This is because, by design, all EB items in $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ have multiple correct interpretations. Results on the full v2 evaluation benchmark are reported for the baseline and best performing models in this work later in Section 7.1.
@@
An important consideration is determining the optimal amount of each synthetic dataset to use alongside $D_{\mathrm{mc}}$ in the training dataset. We adopt an empirical yet systematic approach to address this, yielding a data mix that optimizes $D_{\mathrm{mc}}$ performance while also being highly performant on the synthetic data as much as possible. Details can be found in the appendix. Table 3 presents the final data statistics. Similar to the methodology used for data splits for the MDC  (Narayan-Chen, Jayannavar, and Hockenmaier, 2019), for each of the synthetic datasets, the game logs are split into disjoint test, training, and validation sets such that training target structures do not appear during test or validation. We now have a total of 52419 items in the aggregated training dataset.

### v2

@@ 1 Introduction
The MCBT situates a dialogue task—designed to be asynchronous, asymmetric, and with minimal language constraints—in a simulated 3D Minecraft environment where participants, with constantly shifting perspectives, must refer to and modify a dynamic world. In this task, two players collaborate: an Architect (A) instructs a Builder (B) to construct a target structure using multi-colored blocks (Figure 1). The MDC comprises 509 human-human game logs for this task. Narayan-Chen, Jayannavar, and Hockenmaier (2019) focus on generating Architect utterances, and Jayannavar, Narayan-Chen, and Hockenmaier (2020) focus on end-to-end neural models for building an automated Builder agent. The latter defined the challenging Builder Action Prediction (BAP) task, aimed at predicting the sequence of actions (block placements and/or removals) a human Builder performed at a specific point in a human-human game. Despite the task’s inherent challenges and limited training data, we demonstrate that models trained with a suitable amount of game history, enriched world representations, and a sufficient amount of synthetic, diversified data (via data augmentation techniques) yield promising initial results. We recap relevant prior work on the Minecraft Builder in Section 2. Building on this foundation, we next take a closer look at evaluation and training data for the BAP task, uncovering key challenges and making significant improvements on both fronts to propose BAP v2, an upgraded version of the task. This will allow future work to make more efficient and meaningful progress on it.
@@
Our work on evaluation highlights spatial reasoning as a key bottleneck in model performance, alongside the challenge of limited training data (even with the aforementioned data augmentation). Instead of merely collecting more data (which can be impractical/expensive in realistic scenarios for such complex situated dialogue tasks), we create additional diverse synthetic training data generated from novel Minecraft dialogue and target structure simulators that emulate the MCBT (Section 5). This data, though naturally simpler than it, is rich in spatial relations and referential expressions and we show that it can be successfully used to train better models, even with straightforward training methods (Section 6). We demonstrate that jointly training on combined synthetic and original BAP (augmented) data already improves performance. Further, a straightforward Curriculum Learning method provides an additional boost. Together, the synthetic and augmented BAP data form the BAP v2 training set. Looking ahead, such data could also be crucial for training more sophisticated, data-hungry deep transformer models and training/fine-tuning increasingly large LLMs. We provide a detailed analysis of model performance of our best performing model relative to the baseline in Section 7.
@@
Section 8 provides a detailed discussion on four topics – an additional key challenge in BAP evaluation, the importance of evaluation on the synthetic data, relevant concurrent work, and broader implications of our work that go beyond BAP. The latter further highlights the potential impact of our work and directions for future research. Related work is covered in Section 9. We conclude and discuss future work in Section 10. We plan to release code and data involved in this work.
@@ 2.1.2 MCBT subtasks: Architect Utterance Generation and Builder Action Prediction
We define the Architect Utterance Generation (AUG) Task as the task of generating a suitable Architect utterance at any point in a human-human game at which the human Architect produced an utterance (Narayan-Chen, Jayannavar, and Hockenmaier, 2019), and the Builder Action Prediction (BAP) Task as the corresponding task of generating a suitable Builder action sequence (consisting of block placements/removals) at any point in a human-human game at which the human Builder placed or removed blocks (Jayannavar, Narayan-Chen, and Hockenmaier, 2020). Although both of these tasks consider only a single turn, and assume the context of a game and dialogue between two human players, they are important first steps towards the creation of agents that can complete an entire game in the MCBT with a human counterpart. We note that compared to BAP, Shi, Feng, and Lipani (2022) propose a slightly different building task involving predicting a single action instead of a sequence of actions. Thus, their setup is not directly comparable to ours.
@@ 2.5 Data Augmentation
Since the small size of the training set (3,709 examples) is a major limiting factor for training complex models, in Jayannavar, Narayan-Chen, and Hockenmaier (2020), we generated synthetic variants of the original game logs in the training data by combining three data augmentation techniques: utterance paraphrasing (synonym-based substitutions), color substitution (random color permutations across logs), and spatial transformations (rotating structures and Builder B). Empirically, we found that increasing the training data to 14,836 (4x) items gave the best performance for our baseline model (Section 2.6). We will refer to this as the original BAP (augmented) data.
@@ 2.6 A CNN- and GRU-based baseline model
In Jayannavar, Narayan-Chen, and Hockenmaier (2020), we proposed a neural model for the BAP task. In this work, we build upon it and refer to it as the baseline model. The model (Figure 4) is based on a recurrent encoder-decoder architecture (Sutskever, Vinyals, and Le, 2014; Cho et al., 2014) in which a GRU-based encoder (bottom left box) captures the game context (dialogue and action history) via GloVe embeddings (Pennington, Socher, and Manning, 2014), and a CNN-based encoder (top left box) captures the world state at each time step. The world state encompasses the current built structure, action history and the Builder’s position and orientation. The decoder (right box) predicts one action per time step, based on the game history, the world state at that time, and the last action taken. It consists of another GRU backbone over action sequences (bottom right), and a multi-class classifier that reads in the output of the GRU backbone as well as the world state encoding produced by the CNN to predict either the next action (block placement or removal) to be taken, or a special STOP token that terminates the action sequence. Actions are represented as feature vectors and embedded via feedforward layers before being fed to the decoder GRU. The world state representation gets updated and re-encoded after each predicted action. We train this model with teacher forcing and cross entropy loss on the original BAP (augmented) data, and use greedy decoding (max. sequence length of 10 actions) at test time. This yields an F1 score of 21.2% on the BAP test set.
@@ 3.1 Annotating the test set
We manually annotate items in the test set so that we can identify cases that have to be distinguished for evaluation purposes. Annotators were asked to annotate all items in the same dialog in sequence. For each non-empty board (NEB) item, annotators were first asked whether the context is clear and has a unique interpretation , i.e. clearly specifies what structure to be built next, and there is only a single way of placing that structure on the board (as is typically the case when the board is not empty), or whether the context is unclear, i.e. it is unclear what the Builder should do next, either because more information is required, or no new instruction was given. For empty board (EB) items, annotators could also indicate that the context is clear and has multiple interpretations , i.e. it is clear what structure should be built, even though there are multiple possible placements or orientations within the build region, as is often the case when the board is empty. If the context is clear (with unique or multiple interpretations), a second question asked whether the built structure is correct or incorrect. Details about the annotation setup are provided in the appendix (Section 14). We now illustrate the resulting categories of BAP items with a few examples.
@@ 5.1.2 High-level Algorithm
    Optional Clarification Exchange: Some information may be omitted from the instruction with a small probability, allowing B to ask a clarification question. An appropriate exchange is then generated, with B asking and A providing the missing information, with template-based lexical and syntactic variations. E.g., in Figure 11, the block’s color is omitted, prompting B to ask, "what color?" and A to reply, "red".
@@
We implement this general framework in three distinct ways, resulting in three different simulators. Each simulator uniquely instantiates steps 1, 3, and 4, along with independent target structure generation, while steps 2, 5, and 6 remain largely consistent across simulators. For brevity, we provide only a summary of each simulator in the following sections and provide more details in the appendix (Section 11).
@@ 5.2 Random target structures and dialogs
The dialogs correspond to random target structures generated dynamically during the dialog simulation (Section 5.1.2) which runs for a (bounded) random number of iterations. Blocks are placed/removed randomly, and one at a time in each iteration, with limited planning on A ’s part for block placement. As a result, the target structure is disordered, and A ’s instructions are fairly low-level, addressing only one block at a time. Most of them use spatial relations relative to arbitrary blocks identified by color and location (e.g., "to the right of the red" or "to the right of the bottom-most red block"). Figure 12 shows an example BAP item from this dataset. A identifies the orange block as a reference block and uses a 1D spatial relation "behind" to specify the target yellow block. We provide more details about the simulator in the appendix (Section 11).
@@ 5.3.1 Shape-based structures
We define six unique elementary shapes: rows, diagonals, T-shapes, L-shapes, U-shapes, and planes. These shapes were chosen for their simplicity and occurrence in MDC dialogs and structures. (Formal definitions of the shapes are provided in the appendix (Section 11.4)). We design a simulator that generates target structures by combining multiple randomly sampled elementary shapes into a single composite structure. Sampling is performed with replacement, allowing multiple instances of the same shape type within a structure. For each shape instance, properties such as size, color, orientation, and exact 3D location are randomly sampled within specific constraints (e.g., minimum and maximum size bounds). Some global constraints are also enforced (similar to those for random structures in Section 5.2, with an added constraint for shapes):
@@ 5.3.2 Blocks-based dialogs for shape-based structures
The dialog simulation (Section 5.1.2) takes the target structure as input and continues iteratively until the target structure is built. Construction proceeds one shape at a time. A ’s planning is involved and follows a heuristic that selects one or more blocks to place in each iteration until all shapes are completed. Consequently, A ’s instructions reference a single block or multiple ones collectively, with each shape decomposed into multiple instructions. Most of A ’s instructions include spatial relations relative to the last block placed by B (e.g., "now, place two red blocks behind the last block"). Figure 15 illustrates an example BAP item from this dataset. The first blue square has been built, and the second (yellow) shape is under construction. A is instructing B to continue building the latter by placing three more blocks, using a 1D spatial relation "on top of" relative to the last block. (In the previous step, A used a 2D spatial relation, "diagonally in front of and to the right of".) We provide more details about the simulator in the appendix (Section 11).
@@ 5.3.3 Shape-based dialogs for shape-based structures
The dialog simulation (Section 5.1.2) takes the target structure as input and proceeds shape by shape, with one complete shape placed per iteration until the target is built. A ’s planning employs simple heuristics to determine the order of shapes and select a target block within each shape for defining spatial relations, which also serves as B ’s starting point for building the shape. A ’s instructions, therefore, describe entire shapes rather than individual blocks, resulting in higher-level, abstractive instructions. Since each target structure consists of two shapes, each placed in one go, a 1D/2D/3D spatial relation is only used to specify the position of the initial block of the second shape relative to the last block placed in the first shape. Additionally, A can also use spatial relations that use B as a spatial anchor (Section 5.1.3), for building either shape. B’s low-level planning also uses heuristics to determine the order of blocks within a shape, mimicking human Builders. Multiple orderings are possible, with some being more optimal or natural than others. E.g., 2D shapes like planes can be built either by column or by row. Additionally, B may also adopt a zigzag pattern, alternating the starting side for each row or column. Figure 16 illustrates a BAP item from this dataset. The first shape (a yellow column) has been built, and A is instructing B to build the second shape (a red row) in one go. A uses the spatial relation "going to the left of you" (using B as a spatial anchor) to indicate the direction for the row, along with the 3D spatial relation "one to the left, two underneath, and one block in front of" relative to the last block (the topmost yellow one), to indicate the starting position for the row. We provide more details about the simulator in the appendix (Section 11).
Figure 16: A BAP item from the shape-based dialogs for shape-based structures
@@ 6 Training with Synthetic Data
Our next objective is to demonstrate that the synthetic data can help train much better models for the BAP task, even with straightforward training methods, thereby addressing the challenge of insufficient training data. We first show that just joint training on the aggregation of the synthetic data and the original BAP (augmented) data already yields a better model (Section 6.1). We then go a step further and show that a simple Curriculum Learning method can further boost performance (Section 6.2). The synthetic data together with the original BAP (augmented) data now constitutes the BAP v2 training dataset, completing the BAP v2 framework. We also illustrate the impact of the synthetic data and training methodologies on training a simple LLM (BERT) and transformer-based model (Section 6.3).
@@ 6.1.1 Experimental Setup
Our setup closely follows the one used for the baseline model $M_{\mathrm{mc}}$ when it was trained on $D_{\mathrm{mc}}$ alone (Section 2.6). We retain the same exact model architecture and hyperparameters as in $M_{\mathrm{mc}}$. Training employs teacher forcing with cross-entropy loss, and decoding is performed via greedy decoding, with a maximum sequence length of 20 actions. Further details are provided in the appendix (Section 12.3). We use the fairer F1 metric to report micro F1 scores on the BAP v2 test set for $D_{\mathrm{mc}}$ (Section 4.4) and the test sets of $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$.  [^6] For fairer F1 on $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$, $\mathcal{M}$ consists of all EB items, unlike $D_{\mathrm{mc}}$, where only a subset is considered (Section 4.1). This is because, by design, all EB items in $D_{\mathrm{bs}}$, $D_{\mathrm{ss}}$, $D_{\mathrm{r}}$ have multiple correct interpretations. Results on the full v2 evaluation benchmark are reported for the baseline and best performing models in this work later in Section 7.1.
@@
An important consideration is determining the optimal amount of each synthetic dataset to use alongside $D_{\mathrm{mc}}$ in the training dataset. We adopt an empirical yet systematic approach to address this, yielding a data mix that optimizes $D_{\mathrm{mc}}$ performance while also being highly performant on the synthetic data as much as possible. More details are provided in the appendix (Section 12.2). Table 3 presents the final data statistics. Similar to the methodology used for data splits for the MDC  (Narayan-Chen, Jayannavar, and Hockenmaier, 2019), for each of the synthetic datasets, the game logs are split into disjoint test, training, and validation sets such that training target structures do not appear during test or validation. We now have a total of 52419 items in the aggregated training dataset.